# Impact of the valence band on Rydberg excitons in cuprous oxide quantum wells

### v1

@@ I Introduction
Electron-hole bound states in bulk cuprous oxide are characterized by large binding energies, allowing for high-resolution measurements of the excitonic transitions of yellow and green Rydberg series [1, 2, 3]. This paves the way for excitonic devices, sensors [4, 5] and potentially room-temperature applications [6, 7, 8]. Due to strong spin-orbit coupling, an accurate theoretical description of exciton spectra in bulk Cu2O requires taking into account the complex valence band structure that goes beyond the simple parabolic band approximation. Effects of the band structure on exciton energies have been investigated recently both theoretically and experimentally [9, 10, 11, 12, 13, 14, 15]. These works have significantly advanced our understanding of the exciton properties in bulk Cu2O, in comparison to the pioneering studies of exciton energies and absorption [16, 17, 18].
@@
Similar to other types of semiconductors such as GaAs [19] or CdSe [20, 21], a greater tunability of electronic and optical properties of quasiparticles in Cu2O can be achieved by fabrication of heterostructures [22, 23]. The first films of cuprous oxide sandwiched between sapphire substrates and thus realizing the weak quantum confinement of lower exciton states have already been obtained [24, 25]. When the size of a particular exciton state is comparable to the thickness of the film, the confinement effect dominates. Therefore, for very thin Cu2O films or for Rydberg exciton states, the regime of strong confinement is reached, and the details of the complex band structure play the role of a small perturbation. As a result, the parabolic approximation of the bands in the vicinity of the $\Gamma$-point is a reasonable choice. However, for weak confinement the effects due to features of the band structure are comparable to the quantum-confinement effects. Thus, the full Cu2O band structure has to be explicitly taken into account.
@@
The quantum confinement appearing due to the band offsets at the heterojunction can, in its simplest form, be modeled by the quantum well (QW) potential. Excitons in QWs exhibit a large variety of interesting phenomena and effects [26, 27]. The theory of excitons and their radiative properties in GaAs-based QWs is well developed [19], though for cuprous oxide finite-sized crystals the energy spectrum as well as the relative oscillator strengths have been obtained only very recently [28, 29]. These works significantly differ from a simplified consideration in Refs. [30, 31] by studying the spectrum as a function of the QW width. A dependence of the exciton energy levels on the width of a QW allows one to study a crossover from weak confinement in the bulk to nearly two-dimensional excitons in case of strong confinement. Besides, above the scattering threshold electron-hole resonances occur. Their spectral lines are usually additionally broadened due to coupling to the continuum. For GaAs-based QWs, the energies and linewidths of the electron-hole resonances have been estimated using Feshbach theory [32] in Refs. [33, 34]. The resonant broadenings can also be computed by both the stabilization and the complex-coordinate-rotation method, as it was done for GaAs-based [35] and Cu2O-based QWs [36]. When varying the confinement strength, so-called bound states in the continuum (BICs), i.e., long-lived Rydberg states of the confined electron-hole pairs, appear in the continuum background [37]. The BICs are uncoupled from the continuum of lower quantum-confinement subbands and thus have zero linewidths [38, 39]. All these effects can be successfully studied numerically by expanding the electron-hole wave function over a basis of B-splines [40, 41]. The wave functions in coordinate space, computed using a B-spline representation, clearly illustrate, e.g., the quenching of states in the strong confinement regime, the localized nature of BICs and, furthermore, allow for the estimation of oscillator strengths.
@@
In Refs. [36, 28, 29], the energy levels of excitons and electron-hole resonances in finite-sized cuprous oxide crystals have been calculated. Moreover, these works shed light on the crossover from weak to strong confinement regime as well as the radiative properties of exciton states of different symmetries. However, they took into account only the effect of the dielectric contrast in the crystal and in the substrate [28] which, for thin films of semiconductors, is usually described by the Rytova-Keldysh potential [42, 43]. The nonparabolicity of the valence bands near the $\Gamma$-point has not been considered so far. Taking into account the nonparabolic energy dispersion of the hole leads to an increase in the number of coupled differential equations originating from the complex valence band. This significantly complicates the numerical treatment of the problem. Fortunately, the B-spline expansion of the exciton wave function still allows for an effective numerical solution of this challenging problem.
@@
The aim of this article is to study the impact of the complex valence band on excitons in cuprous oxide QWs. Besides the quantization over the confinement direction, we consider the Luttinger Hamiltonian [44] or, equivalently, the Suzuki-Hensel Hamiltonian [45] for the hole dispersion in the QW together with a purely parabolic dispersion for the electron. Such a hole dispersion significantly complicates the Schrödinger equation for excitons in cuprous oxide QWs [28, 29]. Moreover, it breaks some exact symmetries of the pure parabolic dispersion, in particular, the rotational invariance around the axis normal to the QW plane, making the angular momentum quantum number $m$ to be only an approximate quantum number. Nevertheless, using the B-spline representation of the wave function, we can calculate the matrix elements of all involved operators and solve the eigenvalue problem to obtain the energy levels as a function of the QW width. We study effects of the complex valence band, originating from coupling of the states with different values of $m$. We also derive expressions for the relative oscillator strengths of the excitonic transitions induced by circularly polarized light. We show that the oscillator strengths are related to derivatives of certain projections of the exciton wave function at zero separation between electron and hole.
@@ II Theory and methods
Let $L$ be the width of a thin layer of cuprous oxide sandwiched between another material with larger band gap. Due to band offsets at the heterojunction, this system can be regarded as a QW. Excitons in the QW are created by circularly polarized laser light along the normal direction to the QW plane. In general, the exciton spectra depend on the orientation of the crystal relative to the QW. In this article, we consider a geometry in which the principal axis $[001]$ of the crystal is aligned along the $z$ direction perpendicular to the QW plane. The Hamiltonian including the impact of the valence band is derived in Sec. II.1 and its symmetry properties are discussed in Sec. II.2. The computation of the oscillator strengths and the numerical algorithm for the diagonalization of the Hamiltonian using an appropriate basis set are explained in Secs. II.3 and II.4, respectively.
@@ II.1 Hamiltonian
$+V_{\mathrm{e}}(z_{\mathrm{e}})+V_{\mathrm{h}}(z_{\mathrm{h}})$ (1)
with
@@
$+2\left(\eta_{1}+2\eta_{2}\right)\boldsymbol{p}_{\mathrm{h}}^{2}\left(\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}}\right)-6\gamma_{2}\left(p_{1\mathrm{h}}^{2}\boldsymbol{I}_{1}^{2}+\mathrm{c.p.}\right)$
$-12\eta_{2}\left(p_{1\mathrm{h}}^{2}\boldsymbol{I}_{1}\boldsymbol{S}_{\mathrm{h}1}+\mathrm{c.p.}\right)$
$-12\gamma_{3}\left[\left\{p_{1\mathrm{h}},p_{2\mathrm{h}}\right\}\left\{\boldsymbol{I}_{1},\boldsymbol{I}_{2}\right\}+\mathrm{c.p.}\right]$
$-12\eta_{3}\left[\left\{p_{1{\mathrm{h}}},p_{2{\mathrm{h}}}\right\}\left(\boldsymbol{I}_{1}\boldsymbol{S}_{\mathrm{h}2}+\boldsymbol{I}_{2}\boldsymbol{S}_{\mathrm{h}1}\right)+\mathrm{c.p.}\right]\Big]\,$, (3)
$H_{\mathrm{SO}}$ $=\frac{2}{3}\Delta\left(1+\frac{1}{\hbar^{2}}\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}}\right)\,.$ (4)
Here, $E_{\mathrm{g}}$ is the gap energy and $\gamma_{i},\eta_{i}$ are the Luttinger parameters.
@@
The inclusion of the full set of Luttinger parameters leads to the non-trivial band structure shown in Fig. 2. Using the notation for the symmetry group $O_{\mathrm{h}}$ [46], we can classify the bands as in the bulk crystal. The lowest conduction band has the symmetry $\Gamma_{6}^{+}$, whereas the uppermost valence bands are split into the uppermost $\Gamma_{7}^{+}$ bands associated with the yellow exciton Rydberg series, and the lower $\Gamma_{8}^{+}$ bands associated with the green series. Note that for excitons in the QW we use the same values of the Luttinger parameters as in the bulk. The material parameters of cuprous oxide are given in Table 1. $\boldsymbol{I}$ and $\boldsymbol{S}_{\mathrm{h}}$ are the quasispin with $I=1$ and hole spin with $S_{\mathrm{h}}=1/2$, whose components fulfill the commutation relations of an angular momentum or spin, $\Delta$ is the spin-orbit coupling parameter, $\{a,b\}=(ab+ba)/2$ is the symmetrized product, and c.p. denotes cyclic permutation.
@@
The potentials $V_{\mathrm{e},\mathrm{h}}(z_{\mathrm{e},\mathrm{h}})$ confine the motion of electron and hole across the QW along the $z$ direction. For a finite-sized crystal surrounded by vacuum or air, the confinement potential of the QW can be well approximated by an infinite potential well,
The energies of electron and hole states in such a QW,
@@
can be described by two quantum numbers $N_{\mathrm{e}},N_{\mathrm{h}}=1,2,3,\dots$. These quantum-confinement energies play the role of scattering thresholds, where bound states or resonances can accumulate [36]. Note that we use the same dielectric constant $\varepsilon=7.5$ for the crystal and the substrate and thus do not consider the dielectric contrast described by the Rytova-Keldysh potential [42, 43] for the interaction between electron and hole as discussed, e.g., in Ref. [28]. We also do not consider central-cell corrections, which can affect even-parity excitons [49].
In contrast to the bulk with its translational invariance in three dimensions, for excitons in QWs there is a translational invariance only in the two-dimensional $(x,y)$ QW plane. We introduce relative and center-of-mass coordinates (with $m_{\mathrm{h}}=m_{0}/\gamma_{1}$)
@@
$H=H_{\mathrm{diag}}+H_{\Delta m=0}+H_{\Delta m=1}+H_{\Delta m=2}+H_{\Delta m_{F}=4}$ (10)
with
@@ II.2 Symmetry properties
In comparison to the hydrogenlike model of excitons in a QW [29], the inclusion of the complex valence band leads to a further reduction of the symmetry. We are considering the situation in which the [001] axis of the crystal is perpendicular to the QW plane. In this case, the symmetry of the system is reduced from $D_{\infty\mathrm{h}}$, and the complete Hamiltonian (10) is now only invariant under the symmetry operations of the $D_{4\mathrm{h}}$ symmetry group [46], containing rotations $C_{4}$ and $C_{2}$, inversion $I$, improper rotations $S_{4}$, and reflections $\sigma_{\mathrm{h}}$, $\sigma_{\mathrm{v}}$, and $\sigma_{\mathrm{d}}$, as illustrated in Fig. 3. With $m_{J}=m_{I}+m_{S_{\mathrm{h}}}$ and $m_{F}=m+m_{J}$, the $C_{4}$ symmetry implies that the Hamiltonian only couples basis states with $\Delta m_{F}=0,\pm 4$. Note that this symmetry-derived constraint on matrix elements only applies to the total angular momentum, i.e., individual angular momenta like $m$ or $m_{J}$ are not restricted in this way in their possible couplings, as can be seen by the terms in the Hamiltonian (10).
@@
To understand how these blocks relate to the irreducible representations $\Gamma_{6/7}^{\pm}$, we can look at a state with total angular momentum $F=1/2$. We need to take into account that the eigenstates of the quasispin $I$ do not behave exactly like eigenstates of a standard angular momentum. In $D_{4\mathrm{h}}$, the eigenstates of the quasispin transform like the sum of irreducible representations $\Gamma_{4}^{+}+\Gamma_{5}^{+}=(\Gamma_{2}^{+}+\Gamma_{5}^{+})\times\Gamma_{3}^{+}$, while the eigenstates of a standard spin of 1 behave like $\Gamma_{2}^{+}+\Gamma_{5}^{+}$. Hence, one obtains the total symmetry by the usual addition of angular momenta, followed by a multiplication with $\Gamma_{3}^{+}$. With this prescription in mind, we can use that an angular momentum of $1/2$ transforms according to $\Gamma_{6}^{\pm}$ in $D_{4\mathrm{h}}$ [46]. Multiplying by $\Gamma_{3}^{+}$, we find that a state with total angular momentum $F=1/2$, and thus states with $m_{F}=1/2$ and $m_{F}=-1/2$ transform according to $\Gamma_{7}^{\pm}$.

### v2

@@ I Introduction
Electron-hole bound states in bulk cuprous oxide are characterized by large binding energies, allowing for high-resolution measurements of the excitonic transitions of yellow and green Rydberg series [1, 2, 3]. This paves the way for excitonic devices [4], sensors [5, 6] and potentially room-temperature applications [7, 8, 4]. Due to strong spin-orbit coupling, an accurate theoretical description of exciton spectra in bulk Cu2O requires taking into account the complex valence band structure that goes beyond the simple parabolic band approximation. Effects of the band structure on exciton energies have been investigated recently both theoretically and experimentally [9, 10, 11, 12, 13, 14, 15]. These works have significantly advanced our understanding of the exciton properties in bulk Cu2O, in comparison to the pioneering studies of exciton energies and absorption [16, 17, 18].
@@
Similar to other types of semiconductors such as GaAs [19, 20] or CdSe [21, 22], a greater tunability of electronic and optical properties of quasiparticles in Cu2O can be achieved by fabrication of heterostructures [23, 24]. The first films of cuprous oxide sandwiched between sapphire substrates and thus realizing the weak quantum confinement of lower exciton states have already been obtained [25, 26]. When the size of a particular exciton state is comparable to the thickness of the film, the confinement effect dominates. Therefore, for very thin Cu2O films or for Rydberg exciton states, the regime of strong confinement is reached, and the details of the complex band structure play the role of a small perturbation. As a result, the parabolic approximation of the bands in the vicinity of the $\Gamma$ point is a reasonable choice. However, for weak confinement the effects due to features of the band structure are comparable to the quantum-confinement effects. Thus, the full Cu2O band structure has to be explicitly taken into account.
@@
The quantum confinement appearing due to the band offsets at the heterojunction can, in its simplest form, be modeled by the quantum well (QW) potential [27]. Excitons in QWs exhibit a large variety of interesting phenomena and effects [28, 29]. The theory of excitons and their radiative properties in GaAs-based QWs is well developed [30, 19], though for cuprous oxide finite-sized crystals the energy spectrum as well as the relative oscillator strengths have been obtained only very recently [31, 32]. These works significantly differ from a simplified consideration in Refs. [33, 34] by studying the spectrum as a function of the QW width. A dependence of the exciton energy levels on the width of a QW allows one to study a crossover from weak confinement in the bulk to nearly two-dimensional excitons in case of strong confinement. Besides bound states, above the scattering threshold electron-hole resonances occur. Their spectral lines are usually additionally broadened due to coupling to the continuum. For GaAs-based QWs, the energies and linewidths of the electron-hole resonances have been estimated using Feshbach theory [35] in Refs. [36, 37]. The resonant broadenings can also be computed by both the stabilization and the complex-coordinate-rotation method, as it was done for GaAs-based [38] and Cu2O-based QWs [39]. When varying the confinement strength, so-called bound states in the continuum (BICs), i.e., long-lived Rydberg states of the confined electron-hole pairs, appear in the continuum background [40]. The BICs are uncoupled from the continuum of lower quantum-confinement subbands and thus have zero linewidths [41, 42]. All these effects can be successfully studied numerically by expanding the electron-hole wave function over a basis of B-splines [43, 44]. The wave functions in coordinate space, computed using a B-spline representation, clearly illustrate, e.g., the quenching of states in the strong confinement regime, the localized nature of BICs and, furthermore, allow for the estimation of oscillator strengths.
@@
In Refs. [39, 31, 32], the energy levels of excitons and electron-hole resonances in finite-sized cuprous oxide crystals have been calculated using the two-band model with simple parabolic dispersions of the bands. Although these works shed light on the crossover from the weak to the strong confinement regime as well as on the radiative properties of exciton states of different symmetries, the exciton states were studied within the hydrogenlike model, confined over one dimension. The complex structure of the Cu2O valence subbands near the $\Gamma$ point has not been considered so far. Taking into account the complex valence band dispersion leads to an increase in the number of coupled differential equations corresponding to different projections of the orbital momemtum and spin. This makes the numerical treatment of the problem much more difficult.
The transition from bulk to a monolayer affects exciton states in many ways. Apart from restricting the motion of electron and hole along one dimension, it also leads to a suppression of the exciton kinetic energy, a change of the band structure and a distortion of the Coulomb interaction. The reduction of spatial dimension from three to two dimensions reduces the kinetic energy, thereby increasing the exciton binding energy from $E^{3D}_{b}=\text{Ry}/n^{2}$ to $E^{2D}_{b}=\text{Ry}/(n-1/2)^{2}$, where Ry is the Rydberg energy and $n=1,2,\ldots$ is the principal quantum number [28, 45]. The band structure also depends on the geometry of the crystal structure: the effective mass parameters in the kinetic terms are modified during a crossover from bulk to a monolayer. Moreover, this crossover affects the potential energy of the electron-hole interaction. The dielectric contrast between the thin film and the environment (vacuum or a substrate) distorts the pure Coulomb interaction. For thin films, the resulting screening effect can be described by the Rytova-Keldysh potential [46, 47]. Its impact on exciton states is well known for transition metal dichalcogenide monolayers [48, 49, 50], phosphorene monolayers [51], and has been studied for Dirac materials [52] and cuprous oxide thin films [31].
@@
The aim of this article is to study the impact of the complex valence band on excitons in cuprous oxide QWs. Besides the quantization over the confinement direction, we consider the Luttinger Hamiltonian [53] or, equivalently, the Suzuki-Hensel Hamiltonian [54] for the hole dispersion in the QW together with a purely parabolic dispersion for the electron. Such a hole dispersion significantly complicates the Schrödinger equation for excitons in cuprous oxide QWs [31, 32]. Moreover, it breaks some exact symmetries of the pure parabolic dispersion, in particular, the rotational invariance around the axis normal to the QW plane, causing the angular momentum quantum number $m$ to be only an approximate quantum number. Nevertheless, using the B-spline representation of the wave function, we can calculate the matrix elements of all involved operators and solve the eigenvalue problem to obtain the energy levels as a function of the QW width. We study effects of the complex valence band, originating from coupling of states with different values of $m$. We also derive expressions for the relative oscillator strengths of the excitonic transitions induced by circularly polarized light. We show that the oscillator strengths are related to derivatives of certain projections of the exciton wave function at zero separation between electron and hole.
@@ II Theory and methods
Let $L$ be the width of a thin layer of cuprous oxide sandwiched between another material with larger band gap. Due to band offsets at the heterojunction, this system can be regarded as a QW. Excitons in the QW are created by circularly polarized laser light propagating perpendicularly to the QW plane. In general, the exciton spectra depend on the orientation of the crystal relative to the QW. In this article, we consider a geometry in which the principal axis $[001]$ of the crystal is aligned along the $z$ direction perpendicular to the QW plane. The Hamiltonian including the impact of the valence band is derived in Sec. II.1 and its symmetry properties are discussed in Sec. II.2. The computation of the oscillator strengths and the numerical algorithm for the diagonalization of the Hamiltonian using an appropriate basis set are explained in Secs. II.3 and II.4, respectively.
@@ II.1 Hamiltonian
$+V_{\mathrm{e}}(z_{\mathrm{e}})+V_{\mathrm{h}}(z_{\mathrm{h}}),$ (1)
where
@@
$+2\left(\eta_{1}+2\eta_{2}\right)\boldsymbol{p}_{\mathrm{h}}^{2}\left(\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}}\right)-6\gamma_{2}\left(p_{1\mathrm{h}}^{2}{I}_{1}^{2}+\mathrm{c.p.}\right)$
$-12\eta_{2}\left(p_{1\mathrm{h}}^{2}{I}_{1}{S}_{\mathrm{h}1}+\mathrm{c.p.}\right)$
$-12\gamma_{3}\left[\left\{p_{1\mathrm{h}},p_{2\mathrm{h}}\right\}\left\{{I}_{1},{I}_{2}\right\}+\mathrm{c.p.}\right]$
$-12\eta_{3}\left[\left\{p_{1{\mathrm{h}}},p_{2{\mathrm{h}}}\right\}\left({I}_{1}{S}_{\mathrm{h}2}+{I}_{2}{S}_{\mathrm{h}1}\right)+\mathrm{c.p.}\right]\Big]\,$, (3)
$H_{\mathrm{SO}}$ $=\frac{2}{3}\Delta\left(1+\frac{1}{\hbar^{2}}\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}}\right)\,$, (4)
Here, $E_{\mathrm{g}}$ is the gap energy, $m_{0}$ is the free electron mass, and $\gamma_{i},\eta_{i}$ are the Luttinger parameters.
@@
The material parameters of cuprous oxide are given in Table 1. $\boldsymbol{I}$ and $\boldsymbol{S}_{\mathrm{h}}$ are the quasispin with $I=1$ and the hole spin with $S_{\mathrm{h}}=1/2$, whose components $I_{i}$ and $S_{\mathrm{h}i}$ with $i=1,2,3$ satisfy the commutation relations of an angular momentum or spin, $\Delta$ is the spin-orbit coupling parameter, $\{a,b\}=(ab+ba)/2$ is the symmetrized product, and c.p. denotes cyclic permutation.
@@
The potentials $V_{\mathrm{e},\mathrm{h}}(z_{\mathrm{e},\mathrm{h}})$ confine the motion of electron and hole across the QW along the $z$ direction. For a finite-sized crystal surrounded by vacuum or air, the confinement potential of the QW can be well approximated by an infinite potential well, given in Eq. (7). The energies of electron and hole states in such a QW,
@@
can be described by two quantum numbers $N_{\mathrm{e}},N_{\mathrm{h}}=1,2,3,\dots$. Here, $m_{\mathrm{e}}=0.99m_{0}$ is the electron mass in the crystal and $m_{\mathrm{h}}=m_{0}/\gamma_{1}$ is the hole mass. The quantum-confinement energies (8) play the role of scattering thresholds, where bound states or resonances can accumulate [39].
The inclusion of the full set of Luttinger parameters leads to the non-trivial band structure shown in Fig. 2. Using the notation for the symmetry group $O_{\mathrm{h}}$ [57], we can classify the bands as in the bulk crystal. The lowest conduction band has the symmetry $\Gamma_{6}^{+}$, whereas the uppermost valence bands are split into the uppermost $\Gamma_{7}^{+}$ band associated with the yellow exciton Rydberg series, and the lower $\Gamma_{8}^{+}$ bands associated with the green series. Note that for excitons in the QW we use the same values of the Luttinger parameters as in the bulk.
We use the same dielectric constant $\varepsilon=7.5$ for the crystal and the substrate and thus do not consider the dielectric contrast described by the Rytova-Keldysh potential [46, 47] for the interaction between electron and hole as discussed, e.g., in Ref. [31]. We also do not consider central-cell corrections, which can affect even-parity excitons [58].
In contrast to the bulk with its translational invariance in three dimensions, for excitons in QWs there is a translational invariance only in the two-dimensional $(x,y)$ QW plane. We introduce relative and center-of-mass coordinates
@@
$H=H_{\mathrm{diag}}+H_{\Delta m=0}+H_{\Delta m=1}+H_{\Delta m=2}+H_{\Delta m_{F}=4},$ (12)
where
@@ II.2 Symmetry properties
In comparison to the hydrogenlike model of excitons in a QW [32], the inclusion of the complex valence band leads to a further reduction of the symmetry. We consider the situation in which the [001] axis of the crystal is perpendicular to the QW plane. In this case, the symmetry of the system is reduced from $D_{\infty\mathrm{h}}$, and the complete Hamiltonian (12) is now only invariant under the symmetry operations of the $D_{4\mathrm{h}}$ symmetry group [57], containing rotations $C_{4}$ and $C_{2}$, inversion $I$, improper rotations $S_{4}$, and reflections $\sigma_{\mathrm{h}}$, $\sigma_{\mathrm{v}}$, and $\sigma_{\mathrm{d}}$, as illustrated in Fig. 3. With $m_{J}=m_{I}+m_{S_{\mathrm{h}}}$ and $m_{F}=m+m_{J}$, the $C_{4}$ symmetry implies that the Hamiltonian only couples basis states with $\Delta m_{F}=0,\pm 4$. Note that this symmetry-derived constraint on matrix elements only applies to the total angular momentum, i.e., individual angular momenta like $m$ or $m_{J}$ are not restricted in this way in their possible couplings, as can be seen by the terms in the Hamiltonian (12).
@@
To understand how these blocks relate to the irreducible representations $\Gamma_{6,7}^{\pm}$, we can look at a state with total angular momentum $F=1/2$. We need to take into account that the eigenstates of the quasispin $I$ do not behave exactly like eigenstates of a standard angular momentum. In $D_{4\mathrm{h}}$, the eigenstates of the quasispin transform like the sum of irreducible representations $\Gamma_{4}^{+}+\Gamma_{5}^{+}=(\Gamma_{2}^{+}+\Gamma_{5}^{+})\times\Gamma_{3}^{+}$, while the eigenstates of a standard spin of 1 behave like $\Gamma_{2}^{+}+\Gamma_{5}^{+}$. Hence, one obtains the total symmetry by the usual addition of angular momenta, followed by a multiplication with $\Gamma_{3}^{+}$. With this prescription in mind, we can use that an angular momentum of $1/2$ transforms according to $\Gamma_{6}^{\pm}$ in $D_{4\mathrm{h}}$ [57]. Multiplying by $\Gamma_{3}^{+}$, we find that a state with total angular momentum $F=1/2$, and thus states with $m_{F}=1/2$ and $m_{F}=-1/2$ transform according to $\Gamma_{7}^{\pm}$.